# Effect of generic dark matter halo on transonic accretion onto galactic black holes

### v1

@@ IV Critical point analysis and Global accretion solution
The accretion flow starts subsonically from the outer edge of the accretion disk ($v<C_{s}$) and continues smoothly along a streamline with the infall velocity approaching the speed of light as it reaches the event horizon. Thus, during the course of accretion the flow turns supersonic at a critical radius $r_{c}$ depending on the background geometry, the specific angular momentum and energy of the flow. Hence, the accretion flow is transonic in nature. At the critical radius, both $\mathcal{D}$ and $\mathcal{N}$ in Eq. 30 must vanish identically to maintain the smoothness of the accretion flow. Setting the critical point conditions $\mathcal{D}=\mathcal{N}=0$, we get,
@@
×$\times\left[\frac{1}{r_{c}}\left(\frac{5}{2}+\frac{\lambda^{2}}{r_{c}^{2}-\lambda^{2}}\right)+\frac{f^{\prime}(r_{c})}{2f(r_{c})}\right]^{-1},$ (34)
@@
Using the flow velocity ($v_{c}$) and temperature ($\Theta_{\rm c}$) evaluated at the critical point as initial conditions, we integrate Sections III and 30 from the critical point to the outer edge ($r_{\rm edge}$) and again from the critical point to the horizon ($r_{\rm h}$). Finally, we combine both these segments to get the global transonic accretion solution. If the accretion flow contains only an inner critical point ($r_{\rm in}$), the corresponding flow topology is referred to as I-type, whereas if the flow contains only an outer critical point ($r_{\rm out}$), it is called O-type. Depending on the background geometry and flow parameters, the accretion flow may as well contain both the inner and outer critical points simultaneously. In such cases, if the entropy accretion rate at $r_{\rm in}$ is greater than that at $r_{\rm out}$ (˙${\dot{\cal M}}_{\rm in}>{\dot{\cal M}}_{\rm out}$), the flow is referred to as A-type, whereas for the reversed case, when ˙${\dot{\cal M}}_{\rm in}<{\dot{\cal M}}_{\rm out}$, the flow is called W-type Chakrabarti (1989, 1996). We represent all these types of solutions in Section IV.2. For an A-type solution, the flow coming from the outer edge ($r_{\rm edge}$), connects the horizon ($r_{\rm h}$) only through the outer critical point ($r_{\rm out}$). On the other hand, for a W-type solution, the accretion flow passing through $r_{\rm in}$ only connects the outer edge ($r_{\rm edge}$) and horizon ($r_{\rm h}$).
@@ IV.1 Critical point analysis
We analyze the dependency of the critical points on flow properties $i.e.,$ energy ($\mathcal{E}$) and angular momentum ($\lambda$) in the absence and presence of various dark matter distribution ($M_{\rm halo}=10M_{\rm BH},a_{0}=10^{5}M_{\rm BH}$). In Fig. 3, the surface plot represents the nature of the critical points with $\mathcal{E}$ and $\lambda$. Critical points $r_{c}$ are plotted along the X-axis in the logarithmic scale, $\lambda$ is plotted along the Y-axis and $\mathcal{E}$ is plotted along the Z-axis. The yellow, grey, and green colours denote the saddle, nodal, and spiral-type critical points. The range of critical points is not the same for all DM distributions and exhibits notable deviations from the Schwarzschild case for the HQ-type DM spike. A common feature is that there is an upper and lower limit of energy and angular momentum, beyond which the multiple saddle-type critical points cease to exist. This happens due to the choice of the relativistic equation of state and is in agreement with the findings of previous studies Dihingia et al. (2020). As we increase the angular momentum, the likelihood of forming outer critical points is reduced. Under such conditions, the flow enters the BH only through the inner critical points. Accretion flow with lower specific angular momentum passes through the outer critical points only. As these solutions are adiabatic in nature, the entropy accretion rate is constant throughout the journey from the outer edge to the horizon, passing through a critical point.
@@
To analyse the effect of $M_{\rm halo}$ and $a_{0}$, we refer the reader to Section V, where we focus on the portion of the $M_{\rm halo}-a_{0}$ parameter space for which both the inner and outer critical point exists.
@@ IV.2 Global accretion solution
In this section, we analyze of the variation of flow topologies due to the presence of DM halo. Table 1 summarizes the variation of the flow topologies with specific angular momentum ($\lambda$) for $\mathcal{E}=1.0015$, $M_{\rm halo}=10M_{\rm BH}$, and $a_{0}=10^{5}M_{\rm BH}$ for different types of DM distribution and equal mass Schwarzschild BH. For simplicity, in Fig. 4, we depict the flow topologies around a vacuum Schwarzschild BH and that around a galactic BH of the same mass immersed in an Einasto-type DM halo for the parameters mentioned above. We observe that for the chosen specific flow energy, the specific angular momentum range for the I-type solution for Schwarzschild BH falls within the $\lambda$ range of the I-type solutions for all dark matter profiles. From Fig. 4, we note that the presence of the Einasto-type DM halo changes the flow topology from O-type to A-type for $\lambda=2.875$ and A-type to W-type for $\lambda=3.1$. Increasing the angular momentum as $\lambda=3.36$, the topology of the solution changes from W-type to I-type in the presence of the Einasto-type DM distribution. Thus, changing the DM distribution can, in general, change the topology of the accretion flow depending on the choice of the DM and flow parameters. To elucidate this point further, in the following sections, we study the DM and flow parameter space for the existence of multiple saddle-type critical points.
@@ V Parameter space for multiple critical points
As discussed earlier, the occurrence of multiple (saddle-type) critical points happens only for a specific range of the flow parameters and DM parameters. In this section, fixing the DM parameters, we first scan the $\mathcal{E}-\lambda$ plane for multiple saddle type critical points (Fig.5), then fixing the flow parameters, we repeat the analysis in the $a_{0}-M_{\rm halo}$ plane (Fig. 6). Even within the portion of the parameter space allowing multiple (saddle-type) critical points, the topology of the accretion flow can be different. Though the entropy accretion rate ($\dot{\mathcal{M}}$) remains constant along the flow streamline, $\dot{\mathcal{M}}$ is different for the inner and outer critical points. Thus, we differentiate the accretion flow based on the solution topologies and the entropy accretion rates. In Fig. 5, we show the $\mathcal{E}-\lambda$ parameter space for multiple saddle-type critical points for galactic BHs with different dark matter profiles. In each panel, we portray the corresponding parameter space for a Schwarzschild BH (black dotted line) of the same mass for comparison. The bounded regions with green (solid), blue (dash-dotted), and magenta (dashed) lines stand for three sets of DM parameters ($M_{\rm halo},a_{0}$) = {($10M_{\rm BH},10^{4}M_{\rm BH}$), ($10M_{\rm BH},10^{5}M_{\rm BH}$), ($10^{2}M_{\rm BH},10^{4}M_{\rm BH}$)}, respectively. On the top and left of the boundary of the existence of multiple saddle critical points, the flow topologies are I-type and O-type, respectively. Contrary to the Hernquist, NFW and NFW1 profiles, the accretion flow does not possess multiple saddle type critical points for the HQ-type DM spike distribution with ($M_{\rm halo},a_{0}$)=($10M_{\rm BH},10^{4}M_{\rm BH}$), whereas, for the Einasto type DM distribution the same is also true for ($M_{\rm halo},a_{0}=10^{2}M_{\rm BH},10^{5}M_{\rm BH}$). In  Fig. 5, we observe that compared to a vacuum Schwarzschild BH, the presence of DM halo shifts the allowed region of the parameter space for multiple (saddle) critical points to lower specific energies and smaller specific angular momenta. In cases of Hernquist, NFW, and NFW1 profiles, we observe that for a given halo-compactness ($\Psi=M_{\rm halo}/a_{0}$), increasing the halo mass shifts the upper bound of the specific energy slightly towards higher specific angular momentum. In all cases, we observe that for a fixed halo mass, increasing the halo compactness (lowering $a_{0}$) shrinks the allowed region of the parameter space, which also shifts towards lower specific angular momenta and lower specific energy of the flow. Although the admissible regions of the parameter space for the Hernquist and NFW distributions are almost similar, they are considerably different for the NFW1, Einasto and HQ-type DM spike profiles. The maximum values of $\lambda$ for all the cases considered in Fig. 5 are set by the minimum value of the Keplerian angular momentum (i.e., $\lambda$ at the marginally stable orbit, See Appendix A for details) for the corresponding DM distributions. In each panel, the shaded regions of the allowed parameter space correspond to the W-type solution, whereas the unshaded regions represent the A-type solution. Note that, for the NFW1 profile with $\Psi=0.001$, there exists a small portion of the $\mathcal{E}-\lambda$ parameter space, where only A-type solutions exist.
@@
Figure 5: parameter space in $\mathcal{E}-\lambda$ plane for different dark matter profiles. Here, we take three different ($M_{\rm halo}-a_{0}$) sets. The boundary region represents the parameter space in the $\mathcal{E}-\lambda$ plane for multiple saddle-type critical points. In all the figures, the black dotted line corresponds to a vacuum Schwarzschild black with the same mass. The middle lines (˙$\dot{\cal M}_{\rm in}=\dot{\cal M}_{\rm out}$) in each plot divide the A-type and W-type solutions. See the text for details.
@@
Fig. 6 illustrates the admissible region for multiple critical points in $a_{0}-M_{\rm halo}$ parameter space, corresponding to various sets of flow parameters, namely ($\mathcal{E},\lambda$)={(1.0009,3.0), (1.0009,3.5), (1.0022,3.0)}. The magenta (solid), blue (dashed) and green (dash-dotted) lines represent the lower boundary for the admissible portion of the $a_{0}-M_{\rm halo}$ plane. We observe that for a fixed set of flow parameters, the allowed region of the $a_{0}-M_{\rm halo}$ parameter space may contain either A type, W type or both types of accretion flow. For $\mathcal{E}=1.0022$ and $\lambda=3.0$, the magenta (dotted) line marks the upper bound of the W-type solutions (and lower bound of A-type solutions) for the NFW, NFW1 and Einasto DM distributions. The other solution topologies are marked in the figures respectively.
@@
Table 1: Table showing the specific angular momenta range for different accretion flow topologies around a galactic BH with different types of DM distributions with $M_{\rm halo}=10M_{\rm BH}$, $a_{0}=10^{5}M_{\rm BH}$ and $\mathcal{E}=1.0015$. The maximum value of $\lambda$ for each profile is fixed by the corresponding Keplerian angular momentum $\lambda_{K}$, where, $\lambda_{K}^{\rm Sch}=3.67424$, $\lambda_{K}^{\rm HQ}=3.67460$, $\lambda_{K}^{\rm NFW}=3.6756$, $\lambda_{K}^{\rm NFW1}=3.67519$, $\lambda_{K}^{\rm Ein}=3.67549$ and $\lambda_{K}^{\rm DM-Spk}=3.67625$
@@ VI Solution Properties
In this section, we study how the presence of different DM halo distributions affects the properties of the accretion flow onto the central BH. Following the I-type accretion solution [^2] For other types of solutions, the similar variations also exist., in Fig. 7, we show a representative plot of the flow properties for different halo mass and scale radius for fixed $\mathcal{E}~{}(=1.001)$ and $\lambda~{}(=3.674)$. In the first two columns in Fig. 7, the halo compactness ($\Psi$) increases from $10^{-4}$ to $10^{-2}$ for a fixed $M_{\rm halo}=10^{2}M_{\rm BH}$), whereas in the third column, $M_{\rm halo}$ is increased to $10^{10}M_{\rm BH}$ keeping the halo compactness fixed at $\Psi=0.01$. The first to fifth rows in each column represents the Mach number ($M=v/c_{s}$), temperature ($T$), density ($\rho$), aspect ratio ($H/r$), and optical depth ($\tau_{\rm eff}$), respectively. For comparison in each panel, we also plot the corresponding flow property around a vacuum Schwarzschild BH. We observe that the presence of DM halo shifts the critical point to smaller radii and decreases the flow velocity (as well as Mach number) close to the BH, in general. As the mass accretion rate of the flow is assumed to remain constant throughout, the local mass density varies inversely with the flow velocity (see Eq. 26). So, when the flow velocity (as well as Mach number) is lower than the Schwarzschild BH, the density becomes higher and vice versa. The constant specific energy of the accretion flow with reduced flow velocity results in increased thermal energy of the accreting fluid. This raises the local temperature resulting in an increased thermal pressure, which puffs up the disk, making it quasi-spherical. Thus, at a fixed radius the flow with higher temperature has a higher aspect ratio ($H/r$).
@@
where $\tau_{a}=k_{\rm s}\rho H$ is the optical depth due to Thompson scattering and $\tau_{a}=k^{\rm ff}_{\rm R}\rho H$ is the absorption optical depth. We note from Eqs. 35 and 36 that the optical depth depends on the local flow temperature, density and disk height. Since the presence of DM halo changes each of these flow properties, the exact deviation of the local optical depth is the result of an interplay between them. However, we note from Fig. 7 that the presence of DM halo does not make the accretion disk optically thick. The disk is particularly optically thinner close to the inner edge from where maximum emission occurs (see Section VII for details).
Comparing the first and second columns, we note that increasing the halo compactness increases the deviation from the Schwarzschild BH with the Einasto profile showing the maximum variation. Apart from the HQ-type DM spike profile, increasing the halo mass also increases the deviation of the flow properties from those of the Schwarzschild BH. In the case of HQ-type DM spike distribution, the density profile has a strong nonlinear dependence on the halo mass. For this particular case, it is observed that for small $M_{\rm halo}$ the DM density at a fixed radius decreases with the halo mass, however for large $M_{\rm halo}$, the behaviour is reversed. This change in the halo density with $M_{\rm halo}$ strongly influences the background geometry, which in turn affects the effective potential, causing the flow properties to behave nonlinearly with the halo mass (see Appendix B for details on HQ-type DM spike profile).
@@ VII Radiative emission properties
Figure 8: Plot of the spectral energy distribution (top row) for different DM distributions for different values of the halo mass and scale radius: $M_{\rm halo}=100M_{\rm BH},a_{0}=10^{6}M_{\rm BH}$ (left), $M_{\rm halo}=100M_{\rm BH},a_{0}=10^{4}M_{\rm BH}$ (middle), $M_{\rm halo}=10^{10}M_{\rm BH},a_{0}=10^{12}M_{\rm BH}$ (right). Halo compactness $(M_{\rm halo}/a_{0})$ changes from $10^{-4}$ (left panel) to $10^{-2}$ (middle panel) for a fixed halo mass, whereas it is kept fixed at $0.01$ for right panel ($M_{\rm halo}=10^{10}M_{\rm BH}$) with $M_{\rm BH}=5\times 10^{8}M_{\odot}$. The solid black curve in each plot shows the SED for a Schwarzschild BH of the same mass. The insets show the location of the peak of the SEDs in each case. The flow parameters are $\lambda=3.674$ and $\mathcal{E}=1.001$. See the text for details.
@@
In Fig. 8, we show the spectral energy distribution (SED) of galactic BHs for the different DM distributions with varying halo mass and compactness, ($M_{\rm halo}/M_{\rm BH}$, $\Psi$)={(100, $10^{-4}$), (100, 0.01), ($10^{10}$, 0.01) }. We choose the specific energy and angular momentum of the flow as $\mathcal{E}=1.001$ and $\lambda=3.674$. We choose $\lambda<\lambda_{\rm K}^{\rm Sch}$ to ensure the flow remains sub-keplerian all throughout, where $\lambda_{\rm K}^{\rm Sch}$ is the Keplerian angular momentum for Schwarzschild BH (see Appendix A for details). We observe that the presence of DM causes significant deviation of the SEDs from that of a vacuum Schwarzschild BH of the same mass. The SED also shows a sharp cutoff at $\nu\approx 10^{21}-10^{22}$ Hz, which is governed by the electron temperature at the inner edge of the accretion disk. The presence of a DM halo increases the cutoff frequency from that of the vacuum Schwarzschild BH. Although qualitatively similar, the SEDs for a given DM distribution show significant quantitative variation with the halo mass and compactness. For the low compactness of the DM halo, the spectral energy distribution at high frequency is almost similar to that of the vacuum Schwarzschild BH, whereas the deviation from Schwarzschild BH is relatively large in the low frequencies. On the other hand, for very high compactness, SED shows an even stronger deviation from the vacuum Schwarzshild BH. For a given halo mass, as the halo gets more and more compact, the spectral luminosity increases both in the high and low-frequency regimes. The peak luminosity is also higher for more compact halos and occurs at a relatively lower frequency (except for the Einasto and NFW1 distribution, where it shifts to a higher frequency). The location of the peak of SED corresponds to the frequency and, hence, the energy of the electrons that provide the maximum contribution to the emitted radiation. The cutoff frequency is also higher for more compact halos, which is again an artefact of the higher flow temperature for more compact halos. Though the changes in the SED with halo mass for a given halo compactness are even smaller, we observe the following general trend. The peak luminosity of the Hernquist, NFW, NFW1 and Einasto profiles slightly increases with the halo mass (left to right in Fig. 8) and shifts towards a higher frequency. However, for the HQ-type DM spike profile, the peak spectral luminosity decreases and shifts towards higher frequency as the halo mass is increased. For Hernquist and NFW profiles, the spectral luminosity increases with the halo mass in both low and high-frequency ranges. For NFW1, Einasto and HQ-type DM spike distributions, though the spectral luminosity at high frequencies increases with the halo mass, the behaviour is reversed at low frequencies. The cutoff frequency stays unaffected by the change in halo mass for a given halo compactness except for the HQ-type DM spike distribution, where it increases with the halo mass. Integrating the spectral luminosity over all frequencies, we get the bolometric disk luminosity as,
[…]
@@
Figure 9 shows the variation of the bolometric luminosity with the halo compactness and halo mass for a given value of the flow energy [^3] As the $\mathcal{E}$ is increased keeping a constant $\lambda$ (=3.674), the difference with the Schwarzschild bolometric luminosity starts decreasing. At the value of the specific energy at which the Schwarzschild bolometric luminosity is maximum, all the DM distributions show a reduced bolometric luminosity. and angular momentum. For constant halo mass, as the halo compactness increases, the luminosity for each DM distribution increases from the Schwarzschild value reaching a maximum and then decreasing sharply. The compactness at which the maximum bolometric luminosity is obtained is different for different DM distributions. The maximum value of the luminosity for the Hernquist, NFW, NFW1 and Einasto profiles is almost the same; however, for the DM spike profile, the maximum attained luminosity is substantially smaller. In Fig. 9, we observe that for constant halo mass, as the halo compactness increases, the maximum bolometric luminosity for the Einasto distribution is obtained, followed by the NFW1, Hernquist, NFW and the HQ-type DM spike distributions. As discussed in Section VI, when the compactness of the DM halo increases, the critical point shifts to smaller radii and the temperature of the accretion flow also increases, which in turn increases the luminosity. However, the increasing temperature also puffs up the accretion disk. Since our analysis is valid when the aspect ratio is less than unity ($H/r\leq 1$), the effective portion of the accretion disk contributing to the luminosity reduces, decreasing the luminosity at high compactness.

### v2

@@ IV Critical point analysis and Global accretion solution
The accretion flow starts subsonically from the outer edge of the accretion disk ($v<C_{s}$) and continues smoothly along a streamline with the infall velocity approaching the speed of light as it reaches the event horizon. Thus, during the course of accretion, the flow turns supersonic at a specific radius depending on the background geometry, the specific angular momentum and energy of the flow. Hence, the accretion flow is transonic in nature. At the critical radius ($r_{c}$), both $\mathcal{D}$ and $\mathcal{N}$ in Eq.˜30 must vanish identically to maintain the smoothness of the accretion flow. Setting the critical point conditions $\mathcal{D}=\mathcal{N}=0$, we get,
@@
$\times\left[\frac{1}{r_{c}}\left(\frac{5}{2}+\frac{\lambda^{2}}{r_{c}^{2}-\lambda^{2}}\right)+\frac{f^{\prime}(r_{c})}{2f(r_{c})}\right]^{-1},$ (34)
@@
Using the flow velocity ($v_{c}$) and temperature ($\Theta_{\rm c}$) evaluated at the critical point as initial conditions, we integrate Sections˜III and 30 from the critical point to the outer edge ($r_{\rm edge}$) and again from the critical point to the horizon ($r_{\rm h}$). Finally, we combine both these segments to get the global transonic accretion solution. If the accretion flow contains only an inner critical point ($r_{\rm in}$), the corresponding flow topology is referred to as I-type, whereas if the flow contains only an outer critical point ($r_{\rm out}$), it is called O-type. Depending on the background geometry and flow parameters, the accretion flow may as well contain both the inner and outer critical points simultaneously. In such cases, if the entropy accretion rate at $r_{\rm in}$ is greater than that at $r_{\rm out}$ (${\dot{\cal M}}_{\rm in}>{\dot{\cal M}}_{\rm out}$), the flow is referred to as A-type, whereas for the reversed case, when ${\dot{\cal M}}_{\rm in}<{\dot{\cal M}}_{\rm out}$, the flow is called W-type Chakrabarti (1989, 1996). We represent all these types of solutions in Section˜IV.2. For an A-type solution, the flow coming from the outer edge ($r_{\rm edge}$), connects the horizon ($r_{\rm h}$) only through the outer critical point ($r_{\rm out}$). On the other hand, for a W-type solution, the accretion flow passing through $r_{\rm in}$ only connects the outer edge ($r_{\rm edge}$) and horizon ($r_{\rm h}$).
@@ IV.1 Critical point analysis
We analyze the dependency of the critical points on flow properties $i.e.,$ energy ($\mathcal{E}$) and angular momentum ($\lambda$) in the absence and presence of various dark matter distribution ($M_{\rm halo}=10M_{\rm BH},a_{0}=10^{5}M_{\rm BH}$). In Fig.˜3, the surface plot represents the nature of the critical points with $\mathcal{E}$ and $\lambda$. Critical points $r_{c}$ are plotted along the X-axis in the logarithmic scale, $\lambda$ is plotted along the Y-axis and $\mathcal{E}$ is plotted along the Z-axis. The yellow, grey, and green colours denote the saddle, nodal, and spiral-type critical points. The range of critical points is not the same for all DM distributions and exhibits notable deviations from the Schwarzschild case for the HQ-type DM spike. A common feature is that there are upper and lower limits of energy and angular momentum, beyond which the multiple saddle-type critical points cease to exist. This happens due to the choice of the relativistic equation of state and is in agreement with the findings of previous studies Dihingia et al. (2020). As we increase the angular momentum, the likelihood of forming outer critical points is reduced. Under such conditions, the flow enters the BH only through the inner critical points. Accretion flow with lower specific angular momentum passes through the outer critical points only. As these solutions are adiabatic in nature, the entropy accretion rate is constant throughout the journey from the outer edge to the horizon, passing through a critical point.
@@
To analyse the effect of $M_{\rm halo}$ and $a_{0}$, we refer the reader to Section˜V, where we focus on the portion of the $M_{\rm halo}-a_{0}$ parameter space for which both the inner and outer critical points exist.
@@ IV.2 Global accretion solution
In this section, we analyze the variation of flow topologies due to the presence of DM halo. Table 1 summarizes the variation of the flow topologies with specific angular momentum ($\lambda$) for $\mathcal{E}=1.0015$, $M_{\rm halo}=10M_{\rm BH}$, and $a_{0}=10^{5}M_{\rm BH}$ for different types of DM distribution and equal mass Schwarzschild BH. For simplicity, in Fig.˜4, we depict the flow topologies around a vacuum Schwarzschild BH and that around a galactic BH of the same mass immersed in an Einasto-type DM halo for the parameters mentioned above. We observe that for the chosen specific flow energy, the specific angular momentum range for the I-type solution for the Schwarzschild BH falls within the $\lambda$ range of the I-type solutions for all dark matter profiles. From Fig.˜4, we note that the presence of the Einasto-type DM halo changes the flow topology from O-type to A-type for $\lambda=2.875$ and A-type to W-type for $\lambda=3.1$. Increasing the angular momentum as $\lambda=3.36$, the topology of the solution changes from W-type to I-type in the presence of the Einasto-type DM distribution. Thus, changing the DM distribution can, in general, change the topology of the accretion flow depending on the choice of the DM and flow parameters. To elucidate this point further, in the following sections, we study the DM and flow parameter space for the existence of multiple saddle-type critical points.
@@ V Parameter space for multiple critical points
[…]
@@
Figure 5: parameter space in $\mathcal{E}-\lambda$ plane for different dark matter profiles. Here, we take three different ($M_{\rm halo}-a_{0}$) sets. The boundary region represents the parameter space in the $\mathcal{E}-\lambda$ plane for multiple saddle-type critical points. In all the figures, the black dotted line corresponds to a vacuum Schwarzschild black with the same mass. The middle lines ($\dot{\cal M}_{\rm in}=\dot{\cal M}_{\rm out}$) in each plot divide the A-type and W-type solutions. See the text for details.
@@
Fig.˜6 illustrates the admissible region for multiple critical points in $a_{0}-M_{\rm halo}$ parameter space, corresponding to various sets of flow parameters, namely ($\mathcal{E},\lambda$)={(1.0009,3.0), (1.0009,3.5), (1.0022,3.0)}. The magenta (solid), blue (dashed) and green (dash-dotted) lines represent the lower boundary for the admissible portion of the $a_{0}-M_{\rm halo}$ plane. We observe that for a fixed set of flow parameters, the allowed region of the $a_{0}-M_{\rm halo}$ parameter space may contain either A type, W type or both types of accretion flow. For $\mathcal{E}=1.0022$ and $\lambda=3.0$, the magenta (dotted) line marks the upper bound of the W-type solutions (and lower bound of A-type solutions) for the NFW, NFW1 and Einasto DM distributions. The other solution topologies are marked in the figures respectively.
@@
Table 1: Table showing the specific angular momenta range for different accretion flow topologies around a galactic BH with different types of DM distributions with $M_{\rm halo}=10M_{\rm BH}$, $a_{0}=10^{5}M_{\rm BH}$ and $\mathcal{E}=1.0015$. The maximum value of $\lambda$ for each profile is fixed by the corresponding Keplerian angular momentum $\lambda_{K}$, where, $\lambda_{K}^{\rm Sch}=3.67424$, $\lambda_{K}^{\rm HQ}=3.67460$, $\lambda_{K}^{\rm NFW}=3.6756$, $\lambda_{K}^{\rm NFW1}=3.67519$, $\lambda_{K}^{\rm Ein}=3.67549$ and $\lambda_{K}^{\rm DM-Spk}=3.67625$ .
@@ VI Solution Properties
In this section, we study how the presence of different DM halo distributions affects the properties of the accretion flow onto the central BH. Following the I-type accretion solution [^2] For other types of solutions, the similar variations also exist., in Fig. 7, we show a representative plot of the flow properties for different halo mass and scale radius for fixed $\mathcal{E}~(=1.001)$ and $\lambda~(=3.674)$. In the first two columns in Fig.˜7, the halo compactness ($\Psi$) increases from $10^{-4}$ to $10^{-2}$ for a fixed $M_{\rm halo}=10^{2}M_{\rm BH}$), whereas in the third column, $M_{\rm halo}$ is increased to $10^{10}M_{\rm BH}$ keeping the halo compactness fixed at $\Psi=0.01$. The first to fifth rows in each column represents the Mach number ($M=v/c_{s}$), temperature ($T$), density ($\rho$), aspect ratio ($H/r$), and optical depth ($\tau_{\rm eff}$), respectively. For comparison in each panel, we also plot the corresponding flow property around a vacuum Schwarzschild BH. We observe that the presence of DM halo shifts the critical point to smaller radii and decreases the flow velocity (as well as Mach number) close to the BH, in general. As the mass accretion rate of the flow is assumed to remain constant throughout, the local mass density varies inversely with the flow velocity (see Eq.˜26). So, when the flow velocity (as well as Mach number) is lower than the Schwarzschild BH, the density becomes higher and vice versa. The constant specific energy of the accretion flow with reduced flow velocity results in increased thermal energy of the accreting fluid. This raises the local temperature resulting in an increased thermal pressure, which puffs up the disk, making it quasi-spherical. Thus, at a fixed radius, the flow with higher temperature has a higher aspect ratio ($H/r$).
@@
where $\tau_{a}=k_{\rm s}\rho H$ is the optical depth due to Thompson scattering and $\tau_{a}=k^{\rm ff}_{\rm R}\rho H$ is the absorption optical depth. We note from Eqs.˜35 and 36 that the optical depth depends on the local flow temperature, density and disk height. Since the presence of DM halo changes each of these flow properties, the exact deviation of the local optical depth is the result of an interplay between them. However, we note from Fig.˜7 that the presence of DM halo does not make the accretion disk optically thick. The disk is particularly optically thinner close to the inner edge from where maximum emission occurs (see Section˜VII for details).
Comparing the first and second columns, we note that increasing the halo compactness increases the deviation from the Schwarzschild BH with the Einasto profile showing the maximum variation. Apart from the HQ-type DM spike profile, increasing the halo mass also increases the deviation of the flow properties from those of the Schwarzschild BH. In the case of HQ-type DM spike distribution, the density profile has a strong nonlinear dependence on the halo mass. For this particular case, it is observed that for small $M_{\rm halo}$ the DM density at a fixed radius decreases with the halo mass, however for large $M_{\rm halo}$, the behaviour is reversed. This change in the halo density with $M_{\rm halo}$ strongly influences the background geometry, which in turn affects the effective potential, causing the flow properties to behave nonlinearly with the halo mass (see Appendix˜B for details on HQ-type DM spike profile).
@@ VII Radiative emission properties
Figure 8: Plot of the spectral energy distribution for different DM distributions for different values of the halo mass and scale radius: $M_{\rm halo}=100M_{\rm BH},a_{0}=10^{6}M_{\rm BH}$ (left), $M_{\rm halo}=100M_{\rm BH},a_{0}=10^{4}M_{\rm BH}$ (middle), $M_{\rm halo}=10^{10}M_{\rm BH},a_{0}=10^{12}M_{\rm BH}$ (right). Halo compactness $(M_{\rm halo}/a_{0})$ changes from $10^{-4}$ (left panel) to $10^{-2}$ (middle panel) for a fixed halo mass, whereas it is kept fixed at $0.01$ for right panel ($M_{\rm halo}=10^{10}M_{\rm BH}$) with $M_{\rm BH}=5\times 10^{8}M_{\odot}$. The solid black curve in each plot shows the SED for a Schwarzschild BH of the same mass. The insets show the location of the peak of the SEDs in each case. The flow parameters are $\lambda=3.674$ and $\mathcal{E}=1.001$. See the text for details.
@@
In Fig.˜8, we show the spectral energy distribution (SED) of galactic BHs for the different DM distributions with varying halo mass and compactness, ($M_{\rm halo}/M_{\rm BH}$, $\Psi$)={(100, $10^{-4}$), (100, 0.01), ($10^{10}$, 0.01) }. We choose the specific energy and angular momentum of the flow as $\mathcal{E}=1.001$ and $\lambda=3.674$. We choose $\lambda<\lambda_{\rm K}^{\rm Sch}$ to ensure the flow remains sub-keplerian all throughout, where $\lambda_{\rm K}^{\rm Sch}$ is the Keplerian angular momentum for Schwarzschild BH (see Appendix˜A for details). We observe that the presence of DM causes significant deviation of the SEDs from that of a vacuum Schwarzschild BH of the same mass. The SED also shows a sharp cutoff at $\nu\approx 10^{21}-10^{22}$ Hz, which is governed by the electron temperature at the inner edge of the accretion disk. The presence of a DM halo increases the cutoff frequency from that of the vacuum Schwarzschild BH. Although qualitatively similar, the SEDs for a given DM distribution show significant quantitative variation with the halo mass and compactness. For the low compactness of the DM halo, the spectral energy distribution at high frequency is almost similar to that of the vacuum Schwarzschild BH, whereas the deviation from the Schwarzschild BH is relatively large in the low frequencies. On the other hand, for very high compactness, SED shows an even stronger deviation from the vacuum Schwarzshild BH. For a given halo mass, as the halo gets more and more compact, the spectral luminosity increases both in the high and low-frequency regimes. The peak luminosity is also higher for more compact halos and occurs at a relatively lower frequency (except for the Einasto and NFW1 distributions, where it shifts to a higher frequency). The location of the peak of SED corresponds to the frequency and, hence, the energy of the electrons that provide the maximum contribution to the emitted radiation. The cutoff frequency is also higher for more compact halos, which is again an artefact of the higher flow temperature for more compact halos. Though the changes in the SED with halo mass for a given halo compactness are even smaller, we observe the following general trend. The peak luminosity of the Hernquist, NFW, NFW1 and Einasto profiles slightly increases with the halo mass (left to right in Fig.˜8) and shifts towards a higher frequency. However, for the HQ-type DM spike profile, the peak spectral luminosity decreases and shifts towards higher frequency as the halo mass is increased. For Hernquist and NFW profiles, the spectral luminosity increases with the halo mass in both low and high-frequency ranges. For NFW1, Einasto and HQ-type DM spike distributions, though the spectral luminosity at high frequencies increases with the halo mass, the behaviour is reversed at low frequencies. The cutoff frequency stays unaffected by the change in halo mass for a given halo compactness except for the HQ-type DM spike distribution, where it increases with the halo mass. Integrating the spectral luminosity over all frequencies, we get the bolometric disk luminosity as,
$L=\int_{0}^{\infty}L_{\nu_{o}}d\nu_{o}.$ (40)
@@
Figure 9 shows the variation of the bolometric luminosity with the halo compactness and halo mass for a given value of the flow energy [^3] As the $\mathcal{E}$ is increased keeping a constant $\lambda$ (=3.674), the difference with the Schwarzschild bolometric luminosity starts decreasing. At the value of the specific energy at which the Schwarzschild bolometric luminosity is maximum, all the DM distributions show a reduced bolometric luminosity. and angular momentum. For constant halo mass, as the halo compactness increases, the luminosity for each DM distribution increases from the Schwarzschild value reaching a maximum and then decreasing sharply. The compactness at which the maximum bolometric luminosity is obtained is different for different DM distributions. The maximum value of the luminosity for the Hernquist, NFW, NFW1 and Einasto profiles is almost the same; however, for the DM spike profile, the maximum attained luminosity is substantially smaller. In Fig.˜9, we observe that for constant halo mass, as the halo compactness increases, the maximum bolometric luminosity for the Einasto distribution is obtained, followed by the NFW1, Hernquist, NFW and the HQ-type DM spike distributions. As discussed in Section˜VI, when the compactness of the DM halo increases, the critical point shifts to smaller radii and the temperature of the accretion flow also increases, which in turn increases the luminosity. However, the increasing temperature also puffs up the accretion disk. Since our analysis is valid when the aspect ratio is less than unity ($H/r\leq 1$), the effective portion of the accretion disk contributing to the luminosity reduces, decreasing the luminosity at high compactness.